# “I Want It That Way”: Enabling Interactive Decision Support Using Large Language Models and Constraint Programming

### v1

@@ 6.4.1. Suggestion Evaluation
How the system presented its explanations affected participants’ understanding of system suggestions and which preferences were met. Some participants found that the explanations helped them troubleshoot their preferences: “It was upfront when it couldn’t do something. Right. We asked about to take into consideration weather, and it told me right off, I don’t have that (P6)”. However, P2 reported that text-based explanations made it difficult to comprehend the suggestion compared to when it is visually positioned in the calendar. Several participants (P4, P7, P10) simply missed reading the explanations. To help with understanding their own preferences, some participants (P6, P7, P8) wanted the system to explicitly specify why some preferences were satisfied or not, and the lack of such adequate explanations was sometimes a reason for participants to engage in trial and error. For example, P6 mentioned that “it’s not really telling me why it won’t put on the 6th” and proceeded to investigate the reasoning by interacting with the bot. Receiving the same suggestion without adequate adjustments to the explanation was especially frustrating to some participants (P3, P5, P6, P7). Explaining the suggestion’s impact on attendance was a high priority for some participants (P2, P6): “if I have to schedule with ten people, it’s just helpful to understand who cannot attend. (P2)” Instead of an explanation, P5 suggested the system should ask if the unmet constraint could be relaxed: “if it couldn’t find it on that date, maybe it should have asked, if I’m flexible on the time.” In general, understanding the why behind suggestions helped participants focus their attention on making adjustments to their preferences.
@@ 6.4.3. Other considerations
Leveraging LLMs for incorporating user preferences presented both benefits and challenges. As we already described, LLMs enabled all of the participants to freely interact with the system to express their preferences and questions. The system was especially good at handling misspellings that enhanced user experience: “I’m satisfied to verify my hypothesis that it’s kind of resilient to me being sloppy while typing. So that’s a neat thought about experience. (P1)” However, the system also hallucinated or made errors in some components of the system. The system sometimes generated erroneous Python functions such as the example above where it inverted the inequality sign. In another case, the system generated an erroneous explanation. For example, for a 10 am meeting suggestion, the system explained said that the meeting “is not in the morning.” In both of these error cases, the participants were not given enough information about the underlying system (e.g., generated code, list of constraints) to troubleshoot.
@@ 7. Discussion
We used these preferences to design an interactive system for preference incorporation that can refine suggested meeting times in response to user feedback. The system capitalized on the flexibiliy of LLMs to embed natural language scheduling preferences into Python functions that could then be leveraged by a CP solver to generate refined time suggestions. To validate the technical feasibility of this system for preference incorporation we evaluated it on a quantitative benchmark dataset comprised of results from the diary study. The results showed the viability of this hybrid framework, with the LLM components being able to generate valid python code with high precision and recall with respect to the correct underlying scheduling preference. To evaluate the broader interaction between user’s internal preference construction process and the preference incorporation system we ran a small scale user study. While users found the system intuitive and lauded its flexibility, the study highlgihted key challenges in the design of interactive decision support systems. Namely, users found it difficult to use the system without an understanding of the underlying system and the ability to correct system mistakes. Within the context of meeting scheduling, users also noted that communication by chat could at times be slow and burdensome and recommended features that could allow them to get the feedback and preferences of additional attendees. The remainder of this section elaborates on these important design implications for future interactive decision support systems and discusses limitations of our work.
@@ 7.1.1. System Transparency
Building on this, Lebovitz et al. (2022) have emphasized the significance of interrogation practices that allow users to connect their own knowledge with AI predictions. Other research has explored interventions that aid humans in reasoning about the information available to AI systems (Holstein et al., 2023). These interventions may be valuable in our system, enabling users to learn more about what is feasible and what is not as they interact with the system and rely on it appropriately (Lee and See, 2004). Finally, when the MeetMate system presents multiple time suggestions, incorporating contrastive explanations (Miller, 2019)—explanations that compare different time options—could be a valuable addition. This approach can help users make more informed choices by highlighting the differences, including pros and cons, between various suggestions.
@@ 7.1.2. Human-in-the-loop Error Recovery
Multiple participants in the user study noted inconsistencies between their internal preferences and how the system embedded them (e.g., different time ranges for afternoon), however the current system made it challenging for users to correct implicit assumptions made by the system. This was also validated numerically in the quantitative evaluation where even the current state-of-the-art LLM, GPT-4, was unable to have perfect precision, recall, or safeguard accuracy. Given the potential for errors, interactive decision support systems need intuitive ways for users to diagnose and correct mistakes from the system. This presents a challenging technical problem within the current framework as understanding the underlying representation of user preferences, Python functions, presents a high technical barrier for most users. Future work is needed to bridge this technical gap and provide easy accessible ways to diagnose and correct model errors.
@@ 7.1.3. Reducing Elicitation Time
While the system enabled users to express actionable preferences directly to the system, many users noted the back-and-forth effort of chat to be inefficient and burdensome. Users also voiced frustrations with the system being unaware of scheduling conventions such as lunch times. To help reduce the amount of negotiation time between users and the system to find a meeting time, future systems need to be able to incorporate consistent user preferences and scheduling conventions so that users can focus on fine-tuning suggestions based on new contextual preferences. For instance, this could involve additional form options to capture non-negotiable meeting characteristics like working hours and lunch times as suggested by a number of participants. This could also involve leveraging information about past interactions with the system, including retaining preferences from previous scheduling instances to warm-start the current list of preferences. However, future work is needed to understand how to best leverage preferences across sessions as previous work have shown (scheduling) preferences to be highly transient and contextual (Simon et al., 2008; for double-blind review, 2023).
@@ 7.1.4. Multi-User Preference Construction
Both the diary study and the user study highlighted the limitations of approaching meeting scheduling solely from the organizer’s perspective. Participants noted wanting to query attendees for information such as scheduling preferences, timezone, calendar availability, and feedback on the suggested meeting time. Future work is needed to extend this framework for interactive decision support to a distributed setting where meetings can be scheduled collaboratively between an AI system and multiple attendees. However, such a distributed system presents a number of technical challenges in terms of how to aggregate and weight preferences across users, and how to deal with asynchronous communication between multiple attendees. While previous work has studied this problem in static settings using LLMs (Bakker et al., 2022), future work is needed to adapt these methods to iterative settings such as meeting scheduling.
@@ 7.2. Limitations
While our diary study included participants with a diverse set of roles, they were all employees of the same large technology company. Future work is needed to see if our findings on contextual scheduling preferences generalize beyond the technology sector. The study was also conducted during the summer, which may have limited our ability to capture seasonal variations in scheduling preferences (e.g., those related to the school year or weather). Our user study utilized synthetic scheduling tasks on fake calendar data. While one task required users to use their own scheduling preferences, the manufactured nature of the task might have elicited different preferences than a real scheduling task on the user’s own calendar data. We also highlight the limitations of the scale of the user study. Future work is needed to do larger scale analysis of the system’s usability with personal contextual scheduling tasks.
@@ 8. Conclusion
We belive this paper provides a promising first investigation of a new paradigm for interactive decision support, and highlights a number of important design implications for future systems. Though this paper focused on its application to meeting scheduling, we believe this framework has broader potential in other interactive decision support applications such as personalized desk booking, route planning, and supply chain management. Incorporating more sophisticated optimization tools such as commercial solvers (e.g., gurobi, cplex) is a promising future direction for more complicated decision making settings. Interactive decision support is an inherently cross-disciplinary challenge that requires expertise from optimization, human-computer interaction, and machine learning. We encourage others to continue exploring this challenging and potentially transformative new paradigm for decision support.

### v3

@@ 6.4.1. Suggestion evaluation
How the system presented its explanations affected participants’ understanding of system suggestions and which preferences were met. Some participants found that the explanations helped them troubleshoot their preferences: “It was upfront when it couldn’t do something. Right. We asked to take into consideration weather, and it told me right off, I don’t have that (P6)”. However, P2 reported that text-based explanations made it difficult to comprehend the suggestion compared to when it is visually positioned in the calendar. Several participants (P4, P7, P10) simply missed reading the explanations. To help with understanding their own preferences, some participants (P6, P7, P8) wanted the system to explicitly specify why some preferences were satisfied or not, and the lack of such adequate explanations was sometimes a reason for participants to engage in trial and error. For example, P6 mentioned that “it’s not really telling me why it won’t put [the meeting] on the 6th” and proceeded to investigate the reasoning by interacting with the bot. Receiving the same suggestion without adequate adjustments to the explanation was especially frustrating to some participants (P3, P5, P6, P7). Explaining the suggestion’s impact on attendance was a high priority for some participants (P2, P6): “if I have to schedule with ten people, it’s just helpful to understand who cannot attend. (P2)” Instead of an explanation, P5 suggested the system should ask if the unmet constraint could be relaxed: “if it couldn’t find it on that date, maybe it should have asked, if I’m flexible on the time.” In general, understanding the why behind suggestions helped participants focus their attention on making adjustments to their preferences.
@@ 6.4.3. Other considerations
Leveraging LLMs for incorporating user preferences presented both benefits and challenges. As we already described, LLMs enabled all of the participants to freely interact with the system to express their preferences and questions. The system was especially good at handling misspellings that enhanced user experience: “I’m satisfied to verify my hypothesis that it’s kind of resilient to me being sloppy while typing. So that’s a neat thought about experience. (P1)” However, the system also hallucinated or made errors in the LLM components of the system. Table 2 summarizes the hallucinations encountered during the user study, grouped by the type of error made. Overall we found 14 instances of hallucinations during the user study. The majority of errors occured in the Coder component of the system, with the LLM either implementing constraints incorrectly (e.g., by inverting the inequality sign) or by using a different definition of a term than the user expected (e.g., a different understanding of when the afternoon starts). While we expect the former kind of error can be mitigated with improved prompting or with recent advances in the ability of code generating LLMs, the latter errors emphasizes a need for the system to proactively resolve ambiguity with the user. The Information Checker and Constraint Manager also made a small number of errors while deciding whether a given constraint could be coded, or when selecting an action to take respectively. In all of these error cases, the participants were not given enough information about the underlying system (e.g., generated code, list of constraints) to troubleshoot.
Table 2. Summary of LLM hallucinations during user study. # refers to the number of instances where that type of error occured.
Module Error Type # Example
Coder Logic Error. Python Function implements incorrect code. 6 Input: The meeting needs to start after 11am Output: Code flips inequality sign when checking if meeting occurs after 11am:
def meeting_constraint(candidate_time):
    return candidate_time.start.hour < 11
Definition Error. Misalignment between LLM and user def. of term. 2 Input: I want to meet in the afternoon Output: LLM interprets “meeting is in the afternoon” as the meeting ends after 12pm (user expects it to start after 1pm):
def meeting_constraint(candidate_time):
    return candidate_time.end.hour >= 12
Information Checker Logic Error. Incorrect classification of given constraint. 3 Input: I want to meet during 2-3pm this week because its sunny in Seattle Output: LLM states that given constraint cannot be implemented because the LLM “does not have access to weather data”
Constraint Manager Logic Error. Incorrect action taken given user input. 1 Input: Move the meeting ten minutes after the hour. Output: Constraint Manager changes the priority that Desiree Cain can attend the meeting to one.
Multi-Action Error. Manager only responds to subset of request. 2 Input: I need something the next calendar day. Jennifer and myself are optional for the meeting. Output: Constraint Manager lowers the priority that Jennifer and Dana Nguyen attend the meeting, but ignores next calendar day request.
@@ 7. Discussion
We used these preferences to design an interactive system for preference incorporation that can refine suggested meeting times in response to user feedback. The system capitalized on the flexibiliy of LLMs to embed natural language scheduling preferences into Python functions that could then be leveraged by a CP solver to generate refined time suggestions. To validate the technical feasibility of this system for preference incorporation we evaluated it on a quantitative benchmark dataset comprised of results from the diary study. The results showed the viability of this hybrid framework, with the LLM components being able to generate valid python code with high precision and recall with respect to the correct underlying scheduling preference. To evaluate the broader interaction between user’s internal preference construction process and the preference incorporation system we ran a small scale user study. While users found the system intuitive and lauded its flexibility, the study highlighted key challenges in the design of interactive decision support systems. Namely, users found it difficult to use the system without an understanding of the underlying system and the ability to correct system mistakes. Within the context of meeting scheduling, users also noted that communication by chat could at times be slow and burdensome and recommended features that could allow them to get the feedback and preferences of additional attendees. The remainder of this section elaborates on these important design implications for future interactive decision support systems and discusses limitations of our work.
@@ 7.1.1. System transparency
Building on this, Lebovitz et al. (2022) have emphasized the significance of interrogation practices that allow users to connect their own knowledge with AI predictions. Other research has explored interventions that aid humans in reasoning about the information available to AI systems (Holstein et al., 2023). These interventions may be valuable in our system, enabling users to learn more about what is feasible and what is not as they interact with the system and rely on its recommendations (Lee and See, 2004; Schoeffer et al., 2023). Finally, when the MeetMate system presents multiple time suggestions, incorporating contrastive explanations (Miller, 2019)—explanations that compare different time options—could be a valuable addition. This approach can help users make more informed choices by highlighting the differences, including pros and cons, between various suggestions.
@@ 7.1.2. Human-in-the-loop error recovery
Multiple participants in the user study noted inconsistencies between their internal preferences and how the system embedded them (e.g., different time ranges for afternoon), however the current system made it challenging for users to correct implicit assumptions made by the system. This was also validated numerically in the quantitative evaluation where even the current state-of-the-art LLM, GPT-4, was unable to have perfect precision, recall, or safeguard accuracy. Given the potential for errors, interactive decision support systems need intuitive ways for users to detect, diagnose, and correct mistakes from the system (Schemmer et al., 2023; Schoeffer et al., 2024). This presents a challenging technical problem within the current framework as understanding the underlying representation of user preferences, Python functions, presents a high technical barrier for most users. Future work is needed to bridge this technical gap and provide easy accessible ways to diagnose and correct model error. Relatedly, some higher-skilled users may want deeper access (e.g., the ability to directly edit constraint functions, akin to scripting in Excel) than the current system provides. Determining how to cater to these different levels of expertise is an important open question for future research.
@@ 7.1.3. Reducing elicitation time
While the system enabled users to express actionable preferences directly to the system, many users noted the back-and-forth effort of chat to be inefficient and burdensome. Users also voiced frustrations with the system being unaware of scheduling conventions such as lunch times. To help reduce the amount of negotiation time between users and the system to find a meeting time, future systems need to be able to incorporate consistent user preferences and scheduling conventions so that users can focus on fine-tuning suggestions based on new contextual preferences. For instance, this could involve additional form options to capture non-negotiable meeting characteristics like working hours and lunch times as suggested by a number of participants. This could also involve leveraging information about past interactions with the system, including retaining preferences from previous scheduling instances to warm-start the current list of preferences. However, future work is needed to understand how to best leverage preferences across sessions as previous work have shown (scheduling) preferences to be highly transient and contextual (Simon et al., 2008; Sun et al., 2024).
@@ 7.1.4. Multi-user preference construction
Both the diary study and the user study highlighted the limitations of approaching meeting scheduling solely from the organizer’s perspective. Participants noted wanting to query attendees for information such as scheduling preferences, timezone, calendar availability, and feedback on the suggested meeting time. Future work is needed to extend this framework for interactive decision support to a distributed setting where meetings can be scheduled collaboratively between an AI system and multiple attendees. However, such a distributed system presents a number of technical challenges in terms of how to aggregate and weight preferences across users, and how to deal with asynchronous communication between multiple attendees. While previous work has studied this problem in static settings using LLMs (Bakker et al., 2022), future work is needed to adapt these methods to iterative settings such as meeting scheduling. Another concern in multi-user settings is the management of information leakage between users. For instance, a primary concern is ensuring undisclosed preferences, sensitive features, and private calendar details are not inadvertently shared when negotiating meeting times between multiple users. While there has been some work exploring privacy concerns in leveraging LLMs (Brown et al., 2022), future work is needed to analyze the trade-offs between privacy and efficacy of multi-user meeting scheduling and negotiation.
@@ 7.2. Limitations
While our diary study included participants with a diverse set of roles, they were all employees of the same large technology company. Future work is needed to see if our findings on contextual scheduling preferences generalize beyond the technology sector. The study was also conducted during the summer, which may have limited our ability to capture seasonal variations in scheduling preferences (e.g., those related to the school year or weather). Our user study utilized synthetic scheduling tasks on fake calendar data. While one task required users to use their own scheduling preferences, the manufactured nature of the task might have elicited different preferences than a real scheduling task on the user’s own calendar data. Our decision to explore the MeetMate  as a technology probe limits our ability to draw quantitative conclusions about the usability or effectiveness of the system. We also highlight the limitations of the scale of the user study. Future work is needed to do larger scale analysis of the system’s usability with personal contextual scheduling tasks. We investigated our new framework through the lens of meeting scheduling for information workers, an important but narrow representative task for decision support. We note that other scheduling scenarios (e.g., coordinating calendars in a social group or academic setting) present a different set of implicit assumptions, scheduling preferences, and potential users that may limit the applicability of our findings. Future work is needed to evaluate the framework in a broader class of decision making settings. Furthermore, our study focused on the perspective of the meeting organizer without an active involvement of meeting attendees and consideration of their preferences, and we encourage others to build on our framework and the system to explore distributed settings involving multiple stakeholders.
@@ 8. Conclusion
In this paper, we investigated a new hybrid LLM and optimization framework for interactive decision support. We applied this framework to meeting scheduling, highlighting a number of important design implications for future systems. Our framework could be used in other interactive decision support applications such as personalized desk booking, route planning, and supply chain management. Incorporating more sophisticated optimization tools such as commercial solvers (e.g., Gurobi, CPLEX) is a promising future direction for more complicated decision making settings. Interactive decision support is an inherently cross-disciplinary challenge that requires expertise from optimization, human-computer interaction, and machine learning. We encourage others to continue exploring this challenging new approach for decision support.